OpenStudy (anonymous):

Using the method of Lagrange mulitpliers, optimize the function, f(x,y,z)=x^2+y^2+z^2, subject to x+3y-2z=12.

zepdrix (zepdrix):

\[\large f(x,y,z)=x^2+y^2+z^2\]\[\large g(x,y,z)=x+3y-2z=12\]Mmm so it's been a few weeks since I've done these problems.. trying to remember.. We set them up likeeee, this\[\huge \nabla f=\lambda \cdot \nabla g\]So if we take the partial of both sides with respect to X, we get this equation,\[\large f_x=\lambda \cdot g_x\]\[\large 2x=\lambda\]If we continue with the partials, taking with respect to y, then z, we'll get a system of equations.\[\large 2x=\lambda\]\[\large 2y=3\lambda\]\[\large 2z=-2\lambda\]\[\large 12=x+3y-2z\]

zepdrix (zepdrix):

So we have a system of 4 equations, and 4 unknowns (x, y, z, lambda). There are a few different ways to solve systems. But for this one, I think we can easily solve for x y and z in terms of lambda, and then plug that into the constraint to get a value for lambda.\[\large 2x=\lambda \qquad \rightarrow \qquad x=\frac{1}{2}\lambda\]\[\large 2y=3\lambda \qquad \rightarrow \qquad y=\frac{3}{2}\lambda\]\[\large 2z=-2\lambda \qquad \rightarrow \qquad z=-\lambda\]Plugging these into the constraint gives us,\[\large 12=x+3y-2z \qquad \rightarrow \qquad 12=\left(\frac{1}{2}\lambda\right)+3\left(\frac{3}{2}\lambda\right)-2(-\lambda)\]

OpenStudy (anonymous):

THANK YOU!!

zepdrix (zepdrix):

lol :D can you figure the rest out from there?

OpenStudy (anonymous):

ya all i have to do is solve for (lambda) and then plug it into this-->∇f=λ⋅∇g and solve for x,y, and z? wait how do i solve for x,y and z?

zepdrix (zepdrix):

Ah sorry I'm back :O So did you find your lambda value yet?

OpenStudy (anonymous):

ya it's 12/7

zepdrix (zepdrix):

\[\large x=\frac{1}{2}\lambda \qquad \rightarrow \qquad x=\frac{1}{2}\left(\frac{12}{7}\right)\]This will get you your x value, and so on :D

zepdrix (zepdrix):

After you get an x,y,z.. plug them into f to get a function value. The problem with these constraint problems, is that it can be hard to tell if you found a maximum, minimum or whatever else :P But it looks like it doesn't matter, they simply said "Optimize", they didn't want you to classify it any further it looks like.

zepdrix (zepdrix):

Hopefully we did that correctly :3 do you have a book answer you can compare it to?

OpenStudy (anonymous):

No :( i don't have a book answer but what does it mean by optimize? i don't understand that.. i mean okay so i get the f(x,y,z) value and is that the answer?

zepdrix (zepdrix):

Umm maybe you remember back to calc 1, dealing with optimization. Like... finding the greatest possible area of a box, given a constraint on the perimeter. Orrrrr like.. find the dimensions of a box that minimize the cost to create the box, given a constraint on the volume. In the problem we worked on, we optimized our function when given a certain constraint. We don't actually know if this is a local minimum or maximum of f. There's probably something that could be done to determine that, but we don't need to worry about it. :D

zepdrix (zepdrix):

If you solve for x,y,z I think you get these values, and so we would write our answer as,\[\large f\left(\frac{6}{7},\;\frac{18}{7},\;-\frac{12}{7}\right)=\]And ... then whatever value you get :3

OpenStudy (anonymous):

okay that's what i did haha do you mind helping me understand two more problems from calc 3?

zepdrix (zepdrix):

I can try :3

OpenStudy (anonymous):

kk \[\int\limits_{0}^{4}\int\limits_{(x/2)}^{2}e^ydydx\] sketch the region and interchange the limits of integration

zepdrix (zepdrix):

So our boundaries are, x=0 x=4 and y=x/2 y=2 Let's see if we can sketch that effectively.

zepdrix (zepdrix):